How much water must be added to dilute 10 quarts of a solution that is 18% iodine so that the resulting solution will be 15% iodine?

Answer by stanbon(75887)

How much water must be added to dilute 10 quarts of a solution that is 18% iodine so that the resulting solution will be 15% iodine?
----------------------------------------------
Equation:
iodine + iodine = iodine
0.18*10 + 0*x = 0.15(10+x)
---
Multiply thru by 100 to get:
18*10 + 0 = 15*10+15x
3*10 = 15x
x = 2 quarts (amount of water needed in the mixture)
=====================================================
Cheers,
Stan H.